OpenStudy (anonymous):

what is 3/4 equivalent to?

OpenStudy (anonymous):

0.75

OpenStudy (anonymous):

.75

OpenStudy (anonymous):

\[3/4=(3/4)(25/25)=75/100=0.75\]

OpenStudy (anonymous):

6/8